OpenStudy (anonymous):

What is the solution to the equation 2 1 + 2x = 3 ? answers: x ≈ -0.292 x ≈ -0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (anonymous):

Is it 2(1+2x)=3 or 21 + 2x=3?

OpenStudy (anonymous):

is anything missing in the q?

OpenStudy (anonymous):



OpenStudy (anonymous):

your q has the answer, x = -9

OpenStudy (anonymous):

\[ 2^{1+2 x} = 3\\ (1+ 2 x) \ln 2 = \ln 3\\ 2 x \ln 2 = \ln 3 -ln 2\\ x = \frac{\ln 3 -\ln2}{2\ln 2} \] You can finish it now

OpenStudy (campbell_st):

\[2^{ 1 + 2x} = 3\] take the base 2 log of both sides and using the log law \[\log _{s}(s^n) = n\] so \[1 + 2x = \log _{2} 3\] using change of base... to base e \[1 + 2x = \frac{\ln(3)}{\ln(2)} \] \[2x = \frac{\ln(3)}{\ln(2)} - 1\] \[x = \frac{1}{2}[\frac{\ln(3)}{\ln(2)} - 1]\]

OpenStudy (anonymous):

thanks but im still confused how would u solve the next step?

OpenStudy (anonymous):

x = 0.292

OpenStudy (campbell_st):

pick up a calculator base e logs are ln on a calculator

OpenStudy (anonymous):

ok will do i need to find it tho haha hank you tho campbell and aj and others ur help is greatly appreciated.